OpenStudy (anonymous):

Using the definition of limit to show that lim x approaches -3, where 1-5x is under a surd and is equaling to 4. please help

OpenStudy (anonymous):

you mean the "\[\lim_{x\to a}f(x)=L\] if given any \(\epsilon >0\) there exists a \(\delta\) such that if \(|x-a|<\delta\) then \(|f(x)-L|<\epsilon\) definition?

OpenStudy (anonymous):

yes, what I need to to is how exactly do you solve a problem under a surd using the defination.

OpenStudy (anonymous):

you do these by working backwards that is, write down exactly what you want, and then find the \(\delta\) in terms of \(\epsilon\) that gives it first step is really to go from the general to the specific

OpenStudy (anonymous):

you want \[|\sqrt{1-5x}-4|<\epsilon\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets see, maybe if write down what this means exactly, we can figure out how to achieve it

OpenStudy (anonymous):

I did solve but am not really sure about my approach you know....

OpenStudy (anonymous):

\[-\epsilon<\sqrt{1-5x}-4<\epsilon\] \[4-\epsilon<\sqrt{1-5x}<4+\epsilon\] maybe a good start, although there might be a snappier way in any case now we can get rid of the radical to get to the \(x\)

OpenStudy (anonymous):

looks good square all and get \[16-8\epsilon+\epsilon^2<1-5x<16+8\epsilon+\epsilon^2 \\15-8\epsilon<-5x<15+8\epsilon\\ -3+\frac{8}{3}\epsilon>x>3-\frac{8}{3}\epsilon\]

OpenStudy (anonymous):

typo there in the last line

OpenStudy (anonymous):

\[-3+\frac{8}{3}\epsilon>x>-3-\frac{8}{3}\epsilon\]

OpenStudy (anonymous):

now add 3 and put back in absolute value form \[-\frac{8}{3}\epsilon<x+3<\frac{8}{3}\epsilon\] \[|x+3|<\frac{8}{3}\epsilon\]

OpenStudy (anonymous):

which is exactly what you want, since now you can say "let \(\delta=\frac{8}{3}\epsilon\) " and work the algebra backwards to prove that if this is the case, then \[\sqrt{1-5x}-4|<\epsilon\] although usually you get to be done at that step

OpenStudy (anonymous):

thanks hey, got distracted.

OpenStudy (anonymous):

yw